OpenStudy (anonymous):

(g of f)(x) G(x)= x^2+4x-5 F(x)= 5x-5 help please!!!

OpenStudy (turingtest):

F(x)^{2}+4F(x)-5 =(5x-5)^{2}+4(5x-5)-5 =(25x^{2}-50x+25)+(20x-20)-5 =25x^{2}-30x